# Servo Integrated Nonlinear Model Predictive Control for Overactuated Tiltable-Quadrotors

### v1

@@ V-A Robot Platform
We made a tiltable-quadrotor as shown in Fig. 6 to verify the proposed approach. The robot has 400mm wheelbase and 2.77kg weight, of which the size is suitable for indoor flight. The main modules and their communications are illustrated in Fig. 7. The whole system uses Robot Operating System (ROS) for communication. The main algorithms are running in a Khadas VIM4 onboard computer, which has a 2.2GHz Quad core ARM Cortex-A73 and a 2.0GHz Quad core Cortex-A53 CPU. Then it sends control command to the self-designed flight control unit “Spinal”, which employs a STM32H7 series processor and has been employed in many previous works [9, 10]. Next, “Spinal” sends signals to control actuators. Specifically, we choose a T-Motor F55A PROII 6S 4IN1 ESC with DShot protocal to power four T-Motor AT2814 KV900 motors. Each motor is equipped with a 3-blade 9045 propeller. We select Kondo KRS-3302 as servos.
@@
For state estimation, an IMU is installed on “Spinal”, and its data are filtered in an attitude estimator to obtain roll, pitch angles and attitude velocity. In addition, we use an OptiTrack Motion Capture system to get position and attitude. These measurements are input to an EKF-based state estimator to calculate ${}^{W}\hat{\bm{p}}$, ${}^{W}\hat{\bm{v}}$, ${}^{W}_{G}\hat{\bm{q}}$, and ${}^{G}\hat{\bm{\omega}}$. The servo angle $\hat{\bm{\alpha}}$ is directly measured by the Kondo servo without filtering. These estimated states are sent to NMPC for control.
@@ V-B Parameter Identification
Model predictive control is sensitive to model parameters, hence we conducted several parameter identification experiments. The identified parameters can be separated into four categories: geometry parameters, inertial parameters, rotor parameters, and servo parameters.
(a) Identification Result
(b) Verification in Real Flight
Figure 8: System identification results for the servo model, achieving an average accuracy of 93.4%.
The geometric parameters for the $i$th rotor ${}^{G}\bm{p}_{ri}$ were directly obtained from the CAD model. The mass parameter was weighted using a weighing scale, and the inertial matrix was identified by the bifilar-pendulum method [22]. The rotor parameters $k_{t}$, $k_{q}$, $f_{i,\min}$, $f_{i,\max}$ as well as the mapping between command and thrust were identified by an ATI six-axis force/torque senser installed on a testbed. When identifying rotor parameters, we disassembled one motor from the robot and installed it on the testbed. Comparing this data with the flight performance, we report a 28% lost in thrust due to the aerodynamic interference.
To identify the servo model, we removed propellers for safety, commanded the motor rotating in 60% throttle, and sent a step signal to stimulate the servo to 1.0rad. The data were collected using rosbag, and then the time constant of the servo model was identified by the MATLAB System Identification Toolbox, where we used Prediction Error Minimization as evaluation method. The final result is the average of four servos’ results. Although this no propeller setting may introduce some error compared with the real condition, this accuracy is enough for the real-world flight as shown in Fig. 8. Rotor and servo parameters are listed in Table I.
@@ V-C Tracking Test
The tool acados [23] is utilized to solve the NMPC problem, where we use Partial Condensing HPIPM as QP solver and Explicit Runge-Kutta as integrator. The control parameters are listed in Table I, and the integral term is used only for Z axis.
@@
Parameter Value Parameter Value Parameter Value
@@
$k_{I,xy}$ 0 $k_{I,z}$ 5 $k_{I,qxy}$ 0
$k_{I,qz}$ 0 $f_{d,\rm limit}$ 5 N $\tau_{d,\rm limit}$ 1 $\text{Nm}^{2}$
We conducted three experiments to evaluate the proposed algorithm. Initially, the tiltable-quadrotor took off and then was poked by a stick to test disturbance rejection. Next, we carried out an experiment to track several pose points, aiming to evaluate the performance on physically infeasible trajectory. Finally, the robot was controlled to track a 6-DoF pose trajectory.
V-C1 Takeoff and Disturbance Rejection
Takeoff is the basis for all other experiments. During takeoff, the position command $\bm{p}_{r}=\left[0.0,0.0,1.0\right]^{T}$ was sent to the robot. After takeoff, the robot entered hovering state and was disturbed by a stick. The experiment snapshots and data are shown in Fig. 9 and Fig. 10, respectively.
As can be seen in Fig. 10, the robot can achieve the target height with 13% overshoot, and then the integral term slowly corrects the Z position to the reference. In addition, the drift of X and Y position is smaller than 0.05m. The error for three-axis attitudes without disturbance are lower than 0.05rad. After giving force to the robot, it can recover to the hovering state, demonstrating the robustness of the proposed controller.
@@ V-C2 Set-Pose Tracking
The capability of tracking static points was tested in this experiment, and the reference in each DoF can be seen as a step signal. At the beginning, the pose point $\bm{p}_{r}=\left[0.3,0.2,1.2\right]^{T},\bm{q}_{r}={\rm RPY2Quaternion}\left(0.5,0.0,0.3\right)^{T}$ was sent to the robot. After eight seconds, another pose point $\bm{p}_{r}=\left[-0.3,0.0,1.0\right]^{T},\bm{q}_{r}={\rm RPY2Quaternion}\left(0.5,0.5,-0.3\right)^{T}$ was sent. After another eight seconds, the robot was commanded to return to the starting point. The result is plotted in Fig. 11.
From Fig. 11, the robot can track the position and attitude change simultaneously due to the non-simplification of the nonlinear robot model. The tracking performances for Z position and Yaw attitude are generally better than other axes. In addition, it can be found that X exceeds the reference for about 0.1m, Y is always 0.2m smaller than reference, and Pitch is 0.3rad lower than target. We infer this is due to aerodynamic inference from tilting rotors and model error. Another interesting phenomenon is the oscillation after sending the second point. The second pose point requires longer movement then the first one, making NMPC difficult to find a suitable transition trajectory. This difficulty emphasizes the importance of trajectory planning.
Figure 9: Sequential snapshots demonstrating the takeoff and disturbance rejection phases of the experiment, showcasing the robustness of the controller.
Figure 10: Pose data from the takeoff and disturbance rejection experiment, where vertical dash-dotted lines indicate corresponding moments in Fig. 9.
@@ V-C3 Pose Trajectory Tracking
Finally, a pose trajectory was sent to the robot. Let $\omega=2\pi/20$, then the position $\bm{p}_{r}$ was set as $p_{x}(t)=\cos(\omega t),\ p_{y}(t)=\sin(2\omega t)/2,\ p_{z}(t)=0.3\sin(2\omega t+\pi/2)+1.0$, as well as the attitude $\bm{q}_{r}$ was set from Euler angle as $roll(t)=-\sin(2\omega t)/2,\ pitch(t)=0.5\cos(\omega t),\ yaw(t)=\pi/2\cdot\sin(\omega t+\pi)+\pi/2$. The tracking result is displayed in Fig. 12, and the real-world flight snapshot is presented in Fig. 1.
Figure 12: Pose data from the pose trajectory tracking experiment. The yaw angle is limited in $[-\pi,\pi]$, sometimes resulting the sudden change. The tracking RMSE for all directions are X: 0.153m, Y: 0.169m, Z: 0.020m; Roll: 0.096rad, Pitch: 0.201rad, Yaw: 0.081rad.
From the figure, the tracking performance for Z and Yaw are better than other directions. It can be noted that the pitch tracking is worse than roll and yaw, which may due to aerodynamic disturbance. To sum up, the proposed method is able to track position and attitude trajectories simultaneously, while the modelling error and aerodynamic effect should be further considered to reduce the tracking error.
@@ VI Conclusion
In this article, we proposed an NMPC-based control framework for tiltable-quadrotors. Leveraging a full nonlinear model with servo dynamics, the method directly generated thrust and servo angle as control input. We found in simulation that the inclusion of servo dynamics not only enhanced the control performance but also assisted in the convergence. Finally, the algorithm was verified in real world using a self-made robot.
During flight, the aerodynamic disturbance caused by tilting rotors remains a severe problem. Although integral term is adopted in this article, this way weakens the response speed, presenting the need for more advanced disturbance rejection method. In addition, the influence of servo dynamics on convergence may require further mathematical explanation.

### v2

@@ V-A Robot Platform
We made a tiltable-quadrotor as shown in Fig. 5 to verify the proposed approach, and the basic parameters are listed in Table I. The main modules and their communications are illustrated in Fig. 6, where the Robot Operating System (ROS) is used for communication. The NMPC controller is running in a Khadas VIM4 onboard computer, which has a 2.2 GHz Quad-Core ARM Cortex-A73 and a 2.0 GHz Quad-Core Cortex-A53 CPU. Then it sends control commands to the self-designed flight control unit “Spinal”, which employs an STM32H7 series processor and has been employed in many previous works [8, 9]. Next, “Spinal” sends signals to control actuators. Specifically, we choose a T-Motor F55A PROII 6S 4IN1 ESC with DShot protocol to power four T-Motor AT2814 KV900 motors. Each motor is equipped with a 3-blade 9045 propeller. We select Kondo KRS-3302 as servos.
@@
For state estimation, an IMU is installed on “Spinal”, and its data are filtered in an attitude estimator to obtain roll and pitch angles as well as attitude velocity. In addition, we use an OptiTrack Motion Capture system to get position and attitude. These measurements are input to an EKF-based state estimator to calculate ${}^{W}\hat{\bm{p}}$, ${}^{W}\hat{\bm{v}}$, ${}^{W}_{G}\hat{\bm{q}}$, and ${}^{G}\hat{\bm{\omega}}$. The servo angle $\hat{\bm{\alpha}}$ is directly obtained by the Kondo servo without filtering. These estimated states are sent to NMPC for control.
@@ V-B Parameter Identification
Our parameter identification experiments can be separated into three categories: geometry and inertial parameters, rotor parameters, as well as servo parameters.
The geometric parameters for the $i$th rotor ${}^{G}\bm{p}_{r,i}$ were directly obtained from the CAD model. The mass parameter was weighted using a weighing scale, and the inertial matrix was identified by the bifilar-pendulum method [24].
The rotor parameters $k_{t}$, $k_{q}$, $f_{i,\min}$, $f_{i,\max}$ as well as the mapping between command and thrust were identified by an ATI six-axis force/torque sensor fixed on a testbed. To make the result closer to reality, we 3D-printed a servo structure and mounted it near the rotor with $0^{\circ}$ tilting. Compared to directly fixing the rotor on the force sensor, we observed a 19.53% thrust loss for the same command.
(a) Result for servo identification
(b) Result for rotor identification
Figure 7: For servos, the identified time constant is 0.0859 s, achieving an average accuracy of 93.4%. For rotors, the time constant is 0.0942 s with a 0.35-s dead time, where this large dead time is only observed when starting from zero speed. Thus, the time constants for servo and thrust can be similar.
To identify the servo model, we removed the propellers for safety, commanded the motor to rotate at 60% throttle, and sent a step signal to stimulate the servo to 1 rad ($57.30^{\circ}$). The data were collected using rosbag, and then the time constant of the servo model was identified by the MATLAB System Identification Toolbox, where Prediction Error Minimization is used as the evaluation method. The final result is the average of four servos, with one servo’s result as shown in Fig. 7. Although this no propeller setting may introduce some errors, this accuracy is enough for real-world flight.
@@ V-C Tracking Test
The tool acados [25] is utilized to solve the NMPC problem, where we use Partial Condensing HPIPM as QP solver and Explicit Runge-Kutta as integrator. The control parameters are listed in Table I. We conducted three experiments to evaluate the proposed algorithm. Initially, the tiltable-quadrotor took off and then was poked by a stick to test disturbance rejection. Next, we carried out an experiment to track several pose points, aiming to evaluate the performance of physically infeasible references. Finally, the robot was controlled to track a 6-DoF pose trajectory with different speeds.
@@
Param. Value Param. Value Param. Value
Wheelbase $0.4$ m $m$ $2.773$ kg $\alpha_{\rm limit}$$\pm\pi/2$
$I_{xx}$ $0.0417$ $I_{yy}$ $0.0395$ $I_{zz}$ $0.0707\ \text{kg m}^{2}$
@@
$k_{I,z}$ $5$ $f_{d,\rm limit}$ $5$ N
V-C1 Takeoff, Anti-Disturbance, and Landing
Takeoff is the basis for all other experiments. During takeoff, the position command with the current X-Y position and $p_{r,z}=0.6$ m was sent to the robot. After takeoff, the robot entered a hovering state and was then disturbed by a stick. After that, the robot landed. The data are shown in Fig. 8.
From Fig. 8, the robot can achieve the target height with a 14.83% overshoot, and then the integral term slowly corrects the Z position to the reference within 6 s. After giving disturbances to the robot, it can recover to the hovering state within 1 s, demonstrating the robustness of the proposed controller. After the disturbances ($t>21$ s), the tilting angles of about $10^{\circ}$ still exist, possibly implying that the optimizer has fallen into a local minimum. More analysis can be done in the future.
@@ V-C2 Set-Pose Tracking
The capability of tracking static points was tested, and the reference in each DoF can be seen as a step signal. Initially, the pose point $\bm{p}_{r}[\text{m}]=\left[0.3,0.2,1.2\right]^{T},\ \bm{q}_{r}={\rm RPY2Quat}\left(0.5,0.0,0.3\right)^{T}$ was sent to the robot. After eight seconds, another pose point $\bm{p}_{r}[\text{m}]=\left[-0.3,0.0,1.0\right]^{T},\ \bm{q}_{r}={\rm RPY2Quat}\left(0.5,0.5,-0.3\right)^{T}$ was sent. After another eight seconds, the robot was commanded to the starting point. The result is plotted in Fig. 9. Note that we use radian in code but degree in figures for readability.
From Fig. 9, the robot can track the position and attitude changes simultaneously due to the non-simplification of the nonlinear robot model. The tracking RMSE for all directions are X: 0.072 m, Y: 0.029 m, Z: 0.044 m; Roll: 3.250, Pitch: 2.810, Yaw: 4.342. Some oscillations in Y exist from 16 to 22 s, possibly due to aerodynamic disturbances from environments. In addition, the X position has an error of about 0.08 m from 14 to 22 s, indicating the existence of model error. Overall, we demonstrate the capability of setpoint tracking.
Figure 8: Pose data from the takeoff, anti-disturbance, and landing experiment. A disturbance is given to yaw at about 14.5s (the yellow arrow in the left-upper figure), and then another disturbance is given to X at about 20s (the yellow arrow in the right-upper figure).
@@ V-C3 Pose Trajectory Tracking
Finally, a pose trajectory was sent to the robot. Let $\omega=2\pi/T$, then the position $\bm{p}_{r}$[m] was set as $p_{x}(t)=\cos(\omega t),\ p_{y}(t)=\sin(2\omega t)/2,\ p_{z}(t)=0.3\sin(2\omega t+\pi/2)+1.0$, as well as the attitude $\bm{q}_{r}$ was set from Euler angles as $roll(t)=-\sin(2\omega t)/2,\ pitch(t)=0.5\cos(\omega t),\ yaw(t)=\pi/2\cdot\sin(\omega t+\pi)+\pi/2$. We set $T=20$ s (1x) and $T=10$ s (2x) for one trajectory with different difficulties, where the tracking results are displayed in Fig. 10, and the real-world flight snapshot is presented in Fig. 1.
Figure 10: Pose data for tracking one trajectory with different speeds, where 1x means using 20 s and 2x means using 10 s to finish. The plotted commands are for 2x trajectory. The actual servo angles follow their commands with a slight delay.
The tracking RMSE for 1x trajectory are X: 0.071 m, Y: 0.067 m, Z: 0.018 m; Roll: 4.934, Pitch: 2.023, Yaw: 2.789. RMSE for 2x trajectory are X: 0.103 m, Y: 0.085 m, Z: 0.029 m; Roll: 6.740, Pitch: 1.857, Yaw: 3.622, worse than 1x case. The tracking error in some quick turns (e.g., at 6.5 s for Roll-2x) may be due to model error or physical infeasibility, which can be resolved in the future. Overall, we demonstrate the feasibility of the proposed controller for trajectory tracking.
@@ VI Conclusion
In this article, we proposed an NMPC-based control framework for tiltable-quadrotors. Leveraging a fully nonlinear model with servo dynamics, the method directly generated rotor thrust and servo angle as control inputs. We found in the simulation that the inclusion of servo dynamics not only enhanced the control performance but also assisted in the optimization convergence. Finally, the algorithm was verified in the real world using a self-made robot.
In the future, we plan to implement a disturbance observer within the system to achieve offset-free tracking. Additionally, the proposed controller can be generalized to other aerial robots with tilting structures. We also prepare to extend our approach to handle physical interactions.